ABSTRACT. Let C be a closed convex subset of the Hilbert space H, A a bounded linear operator from H to a finite-dimensional Hilbert space Y, b ∈ Y, and K(b) = C ∩A^{−1}(b). A survey is given of recent results concerned mainly with characterizing and computing best approximations of elements x in H from the set K(b).